OpenStudy (anonymous):

Find the GCF of 12, 18, and 84 A. 6 B. 12 C. 216 D. 252

OpenStudy (anonymous):

D. 252

OpenStudy (campbell_st):

6

OpenStudy (campbell_st):

2 x 6 = 12 3 x 6 = 18 14 x 6 = 84

OpenStudy (anonymous):

Go for A...

OpenStudy (anonymous):

6 12/6=2 18/6=3 84/6=14

OpenStudy (anonymous):

Okay I see what you mean